# Distributional Scaling of Emergent Capabilities

### v3

@@ Abstract
This paper explores the nature of sudden breakthroughs in language model performance at scale, which stand in contrast to smooth improvements governed by scaling laws. While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds. In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. Furthermore, we provide a case study of inverse scaling and show that even as the probability of a successful run declines, the average performance of a successful run continues to increase monotonically. We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations. These insights emphasize the role of random variation in the effect of scale on LM capabilities.
@@ 1 Introduction
On most benchmarks, language model (LM) performance is determined by a scaling law [12, 35, 21] that varies smoothly with parameter size and overall training compute. There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks [41].
@@
These sudden breakthroughs fuel one of the most heated debates in modern machine learning. On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LM to acquire specific concepts that permit out-of-distribution generalization [45]. On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent [36]. Here, we argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
@@
We posit that a breakthrough capability is often distinguished not by deterministic responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics. Such effects are currently undetected because scaling laws are usually determined by measuring a single training run at each scale, and resources are rarely committed to correct for random variation by reusing the same hyperparameter settings with different random seeds. Although random variation may be benign when model performance is measured in-distribution [18], previous work suggests that out-of-distribution performance may vary widely across training runs [49, 51], even at larger scales [26]. Since training numerous seeds becomes prohibitively expensive at large scales, we first study length generalization in small models trained on synthetic tasks. In similar settings, existing publications consider only a few training runs and report their summary statistics. By contrast, we characterize their full distribution, demonstrating that these breakthrough tasks exhibit multimodal variation. Our findings can be summarized as follows:
@@
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters. On the counting task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
@@
  • Breakthrough scaling results from bimodal performance distributions. In language modeling and synthetic length generalization case studies, we demonstrate that some random seeds produce linear scale trends while others are emergent. This variation is caused by the bimodal distribution of each skill across random seeds, a property that materializes around breakthrough thresholds in model size. At these scales, “emergence” is a stochastic property. We confirm these capabilities remain bimodally distributed under continuous metrics.
@@
  • While a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of learning that skill may be responding continuously to scale. Modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful model (Section 3). Measuring the Wassterstein distances between distributions at each scale (Section 3.3), we show that bimodality appears abruptly at some minimum capacity scale, but these sudden changes do not necessarily align with emergent trends for individual model runs.
@@
  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood. We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 3.4), a phenomenon also observed in LMs [28]. We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
@@
  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale. In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LMs. Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
@@
Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length $40$ sequences when independently scaling 2(a) width and 2(b) depth.
2 Experiments
2.1 Synthetic Tasks
We measure models’ length generalization [10, 20, 23, 15], one of many compositional properties that lead to emergence [41, 25, 3]. Experimental setup is described below, with details in Appendix A.
Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) [42]. To observe the random performance distribution at each scale, we train our models from hundreds of seeds. We choose model sizes by separately increasing the width and depth hyperparameters, fixing the other hyperparameter. We scale width by increasing the number of 64-parameter heads per layer.
Task: We consider two algorithmic tasks previously studied in Zhou et al. [49]: counting and reverse order addition. Performance on these task ranges widely across random seeds. At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training. Zhou et al. [49] showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
  • Reverse Order Addition with Index Hints: Zhou et al. [51] showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards. Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length. In this work, we refer to this task as the ‘addition task’ despite these modifications.
Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work [17, 51]. The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for addition. Length generalization is then tested at length 60 for count and 40 for addition.
2.2 Natural Language Tasks
We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task [41, 40]. (A second task, CommonSenseQA, is in Appendix E.) We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs. Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences. We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds. During pretraining, a large diverse corpus encourages models to acquire various capabilities. While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete [29]. This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
Task: We consider finetuning LMs for the MMLU benchmark [11], a multiple-choice question-answering task. Achieving strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format. It is this latter ability which leads to emergent trends [41, 13].
Model: We use the Qwen2.5 family of base models [47]. To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head. We perform full-parameter continued pretraining on these reinitialized models.
Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 [34] with the official MMLU auxiliary training data. The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering. We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 20 reinitializations on each data mixture ratio. We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set. We use a learning rate of 1e-5 with linear decay scheduler.
3 Synthetic Experiment Results
@@
Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs. The literature suggests that these emergent behaviors are unlocked at certain model scales [45], implying that scaling curves for different model runs would generally perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly. We will introduce these concepts and our framework through a case study of the addition task from Section 2.1.
@@ 3.1 Emergence is a Sign of Bimodal Variation
First, we report benchmark performance across scales on our synthetic tasks by emulating the conventional approach in reporting results for single seeds. Following Srivastava et al. [41], we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics. As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1, fixing the random initialization and data order seed across scales. [^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales. As seen in Figure 1 and Appendix Figure 8, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
This variation is explained by the bimodality of model performance distributions when varying seed. The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the addition task, performance clusters into high and low component modes at many parameter sizes. In other words, these scales exhibit distinctly bimodal distributions in length generalization capabilities. This variability is precisely what causes some model runs to appear as breakthroughs while others follow a more linear progression. Specifically, when different runs are clustered into very high or very low performance due to their bimodality, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence.
@@ 3.2 Sudden Jumps From Gradual Distribution Shifts
When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although this artifact which will also produce discontinuity in many single-seed scaling curves. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Figure 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
@@ 3.4 Count: An Instance of U-shaped Scaling
We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 9) but ultimately yields a very different scaling effect: a U-shaped curve. Figure 5 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed. This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Figure 4(b).
U-shaped scaling has been observed in LMs, but its causes are not currently well-understood. When the Inverse Scaling Prize [28] solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al. [46] revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Figure 5 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 5: Summary statistics for length generalization (count task). Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently. The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 5; however, the mean of successful runs (defined as >50% EM accuracy) 5 still exhibits continuous improvement.
3.5 Is Bimodality a Mirage?
Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior [36]; conversely, continuous metrics can make apparently emergent curves into smooth curves [41]. We must be particularly cautious about claiming emergence when requiring model outputs to exactly match a target string, as we do for both synthetic tasks. Are our case studies artifacts of thresholding effects? In other words, do their bimodal distributions become unimodal under continuous metrics? To assess the role of the exact match metric in creating apparent bimodality, we instead consider a continuous equivalent to the exact match metric: the minimum probability assigned to any individual token. If the model fails to assign its highest probability to the correct token for any position in the sequence, it receives a score of zero under the exact match metric, even if the correct token was still highly ranked. Likewise, if the model successfully assigns maximum probability to the correct token, exact match considers it to be accurate whether that probability was high or low. We now consider a continuous version of EM accuracy that reports the worst mistake made by the model. This continuous error score for model $f$ on dataset $X$ is based on the per-token loss $L(f(x_{0\ldots i-1}),x_{i})$:
(a) Fixing network depth to 6 layers while scaling up width.
(b) Fixing network width to 8 heads per layer while scaling up depth.
@@ 3.5 Is Bimodality a Mirage?
4 Natural Language Results
After connecting emergence with bimodality in small synthetic settings, we turn to LMs. We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format [41, 13]. To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them. Because Qwen base models perform well for their scale (39.5% for Qwen2.5-0.5B), these experiments effectively require models to re-discover MMLU capabilities after their removal. Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.

### v4

# Random Scaling of Emergent Capabilities

@@ Abstract
Language models famously improve under a smooth scaling law, but some specific capabilities exhibit sudden breakthroughs in performance. While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds. In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. In a case study of inverse scaling, we show that even as the probability of a successful run declines, the average performance of a successful run increases monotonically. We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations. Our observations hold true even under continuous loss metrics, confirming that random variation must be considered when predicting a model’s performance from its scale.
@@ 1 Introduction
On most benchmarks, language model (LM) performance is determined by a scaling law (Hestness et al., 2017; Rosenfeld et al., 2019; Kaplan et al., 2020) that varies smoothly with parameter size and overall training compute. There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks (Srivastava et al., 2023). These sudden breakthroughs fuel one of the most heated debates in modern AI.
@@
On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales provide the capacity to learn specific concepts (Wei et al., 2022). On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent (Schaeffer et al., 2024). We will argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale. In other words, the discontinuities are real—each model firmly either knows or does not know a given concept—but emergent breakthroughs do not always reflect some fixed threshold which permits the concept to be learned. Instead, models may learn the concept at various scales, albeit with changing probability.
@@
We posit that a breakthrough capability is distinguished not by deterministic responses to scale, but by multimodal random variation. In other words, independent training runs cluster in their performance metrics. This observation is undocumented because scaling laws usually plot a single training run at each scale, rarely testing multiple seeds. Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a; b; Qin et al., 2024; Juneja et al., 2023; Li et al., 2025), even at larger scales (Madaan et al., 2024).
By connecting breakthrough scaling with random variation, we challenge the narratives of both the emergence and mirage camps. First, our results contest the position of the mirage or “loss-to-downstream” camp, which claims that effects of scale are predictable and that continuous metrics will smooth out apparently discrete concept learning. We discover clustered multimodal distributions of capabilities, confirming that models unpredictably learn critical discrete concepts—and these concepts are observable even using continuous loss metrics. Furthermore, we complicate the narrative of the competing emergence camp, which often treats specific model sizes as distinct in their capacity. While we confirm that algorithms require some minimum model capacity, that capacity may not be the enormous scale at which emergence is observed. Instead, discontinuous performance jumps are sampled from a continuously changing multimodal distribution, where the “success” component ultimately dominates at larger scales. To extrapolate downstream metrics to large scales, we must predict the likelihood of successful emergence as well as the performance of a successful run.
Since training numerous seeds is prohibitively expensive at large scales, we study partially reinitialized LLMs and toy models. Whereas prior work reports summary statistics across only a few training runs, we characterize the full multimodal performance distribution. We find:
@@
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters. On the count task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
@@
  • Breakthroughs result from bimodal performance distributions. In multiple choice (Section 3.1) and synthetic task length generalization (Section 2.1), some random seeds produce linear scale trends while others are emergent. This variation is caused by the bimodal distribution of each skill across random seeds (Section 2.2), a property that materializes around breakthrough thresholds in model size. At these scales, emergence is a stochastic property (Section 2.3).
  • Bimodal variation persists under continuous metrics. In the emergence debate, one main position (Schaeffer et al., 2024) is that breakthroughs are caused by measuring discontinuous metrics such as accuracy rather than loss. For our tasks, continuous loss metrics can remain visibly bimodally distributed, particularly when there is a more even split of failed and successful runs. We confirm these loss distributions on both synthetic (Section 2.4) and natural (Section 3.4) tasks.
@@
  • When a scale curve exhibits sudden discontinuous improvement in a skill, the probability of learning that skill may be changing continuously. Treating the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful run (Section 2.2). Although bimodality can appear abruptly at a minimum capacity scale (Section 2.5), these sudden distributional changes do not necessarily align with breakthrough scales for individual model runs.
@@
  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood. We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 2.6), a phenomenon also observed in LMs (McKenzie et al., 2022). We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
@@
  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale. In Section 3, we extend our analysis to the task of multiple choice question answering by training partially reinitialized LMs. Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
@@
Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length 4040 sequences when independently scaling 2(a) width and 2(b) depth.
2 Synthetic Length Generalization Experiments
@@
Usually, performance breakthroughs are measured on a single model per scale or, at most, the average of a few runs. The literature suggests that these emergent capabilities unlock at specific model scales (Wei et al., 2022), implying that scaling curves for different seeds would perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that performance is sampled from a stochastic distribution which changes gradually even as individual scaling curves jump abruptly.
2.1 Tasks and Setup
After training models on synthetic tasks, we measure their length generalization (Graves et al., 2016; Kaiser and Sutskever, 2015; Lake and Baroni, 2018; Hupkes et al., 2020), one of many compositional properties that can lead to conceptual breakthroughs (Srivastava et al., 2023; Löwe et al., 2024; Chen et al., 2024). Further experimental details are in Appendix A.
Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) (Su et al., 2024). To observe the random performance distribution at each scale, we train our models from hundreds of seeds. We choose model sizes by separately adjusting either the width (number of 64-parameter heads per layer) or depth hyperparameter.
Task: We consider two algorithmic tasks previously studied in Zhou et al. (2024a): counting and reverse order addition. Performance on these task ranges widely across random seeds. At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training. Zhou et al. (2024a) showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
  • Addition: Zhou et al. (2024b) showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards. Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length. We use this modified form of reverse order addition with index hints as our addition task.
Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following prior work (Jelassi et al., 2024; Zhou et al., 2024b). The lengths of examples are sampled uniformly from 1 to the maximum training length (30 for count and 35 for addition). Length generalization is then tested at length 60 for count and 40 for addition.
@@ 2.2 Emergence is a Sign of Bimodal Variation
What do the scaling curves for length generalization tasks look like when we generate them in conventional ways? Following Srivastava et al. (2023), we take the vector of model performances for the count task at test length 60 across different scales with a fixed initialization and shuffle seed. [^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales. We calculate the task’s breakthroughness and linearity metrics on each seed. As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1. As seen in Figure 1 and Appendix Figure 10, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
This variation is explained by the bimodality of model performance distributions when varying seed. Figure 2 illustrates that, for a population of models independently trained on the addition task, length generalization clusters into high and low component modes at many parameter sizes. This clustering produces distinctly bimodal performance distributions, causing some model runs to appear as breakthroughs while others follow a more linear progression. When bimodally distributed runs cluster into very high and very low performance components, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence. Furthermore, these differences can be a source of oscillating scale curves, as Srivastava et al. (2023) documented in their lowest-linearity tasks.
@@ 2.3 Sudden Jumps From Gradual Distribution Shifts
When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although the underlying bimodality also produces discontinuity in many single-seed scaling curves. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Figure 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
@@
2.4 Is Bimodality a Mirage?
(a) Fixing network depth to 6 layers while scaling up width.
(b) Fixing network width to 8 heads per layer while scaling up depth.
Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior (Schaeffer et al., 2024); conversely, continuous metrics turn apparent emergence into smooth curves (Srivastava et al., 2023). We must be particularly cautious about claiming emergence when requiring outputs to exactly match a target string, as we have so far. Are our case studies artifacts of thresholding effects? Do our bimodal distributions become unimodal under continuous metrics?
To avoid thresholding artifacts, we use a continuous equivalent to the exact match metric: the maximum loss assigned to any individual token. Each token is individually computed in the correct output context, so the entire sequence represents a fixed set of continuous per-token loss functions. Because the maximum of a fixed set of Lipschitz-continuous functions is always Lipschitz-continuous, this error score is guaranteed to be continuous as long as per-token loss is continuous. For model $f$ on dataset $X$, the continuous error score based on per-token loss $L(f(x_{0\ldots i-1}),x_{i})$ is:
2.5 Bimodality Emerges Abruptly
@@ 2.6 Count: An Instance of U-shaped Scaling
We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 11) but yields a very different scaling effect: a U-shaped curve. Figure 6 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed. This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Figure 5(b).
U-shaped scaling has been observed in LMs, but its causes are not currently well-understood. When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreasing with scale—for large models, Wei et al. (2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Figure 6 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 6: Summary statistics for length generalization (count task). Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently. The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 6; however, the mean of successful runs (defined as >50% EM accuracy) 6 still improves continuously.
@@
3 LLM Experiments
After connecting emergence with bimodality in small synthetic settings, we turn to LMs. We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format (Srivastava et al., 2023; Hu and Frank, 2024). To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them. Because Qwen base models perform well for their scale (39.5% for Qwen2.5-0.5B), these experiments effectively require models to re-discover MMLU capabilities after their removal. Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.
3.1 Data and Setup
We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task (Srivastava et al., 2023; Snell et al., 2024). We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs. Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
Our procedure can be regarded as continued pretraining: we mimic the model’s original pretraining objective with a dataset of English language sequences. We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds. During pretraining, a large diverse corpus encourages models to acquire various capabilities. While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete (Merrill et al., 2023). This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
Task: We test LMs on MMLU Hendrycks et al. (2021), a multiple-choice question-answering benchmark. Strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format. It is this latter ability which leads to emergent trends (Srivastava et al., 2023; Hu and Frank, 2024).
Model: We use the Qwen2.5 family of base models Yang et al. (2024). To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head. We perform full-parameter continued pretraining on these reinitialized models.
Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 Raffel et al. (2023) with the official MMLU auxiliary training data. The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering. We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 80 reinitializations on each data mixture ratio. We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set. We use a learning rate of 1e-5 with linear decay scheduler.
We focus our analysis on experiment results from Qwen2.5-0.5B. In Appendix E, we include results from Qwen2.5-1.5B (Appendix 17) as well as a second task CommonSenseQA (Talmor et al., 2019) (Appendix E.2).